OpenStudy (anonymous):

can someone help please Jose is paying for a six dollar meal using bills in his wallet. He has four one dollar bills, three five dollar bills, and two ten dollar bills. If he selects two bills at random, one at a time from his wallet, what is the probability that he will choose a one dollar bill and a five dollar bill to pay for the meal? Show your work.

OpenStudy (anonymous):

@LegendarySadist

OpenStudy (anonymous):

can you guys help

OpenStudy (anonymous):

I'm here to see how someone else solves it - Probability was never my best subject

OpenStudy (anonymous):

agreed

OpenStudy (anonymous):

i could be totally wrong but ill take a guess

OpenStudy (anonymous):

i know this math subject sucks

OpenStudy (anonymous):

hate probability like i do not get my teacher explains and i am like uhhhh

OpenStudy (anonymous):

total is 4 + 3 + 2 = 9, first pull he has 4 / 9 chance to pull one dollar bill now total bills is 8 second pull he has 3/8 chance to to pull a five dollar bill so 4/9 * 3/8 = answer?

OpenStudy (anonymous):

its competition who will win

OpenStudy (anonymous):

that might be wrong i dont know

OpenStudy (anonymous):

But what if he gets a five dollar bill then a one dollar?

OpenStudy (anonymous):

^^

OpenStudy (anonymous):

seems right i know you have to add first

OpenStudy (anonymous):

How many bills does he have in total? 4 + 2 + 2 = 8 The probability he picks a $10 bill first is therefore (2/8) = (1/4) The probability he picks a $10 bill second, given he didn't pick one first, is (2/7) because there are still two $10 bills in his wallet but only seven bills left Similarly, the probability he picks a $1 bill first is (4/8) = (1/2) The probability he picks a $1 bill second, given he didn't pick one first, is (4/7) because there are still four $1 bills in his wallet but only seven bills left So the probability he picks exactly $11 out with a $1 and $10 bill is the probability either he picks $10 then $1, or $1 then $10, which is: (1/4) * (4/7) + (1/2) * (2/7) = (1/7) + (1/7) = 2/7

OpenStudy (anonymous):

Isn't it 4+3+2=9?

OpenStudy (anonymous):

@nicxle.62 thats not what the question is asking

OpenStudy (anonymous):

uhhh okay um i think @LegendarySadist is right

OpenStudy (anonymous):

beause you do have to add

OpenStudy (anonymous):

woah woah, I'm not saying that's the solution

OpenStudy (anonymous):

no i know you have to add first then do the rest ya i knows but i suck at this

OpenStudy (anonymous):

So I think I have the question figured out.

OpenStudy (anonymous):

i know how to add this because i saw this in other questions like this but different XD

OpenStudy (anonymous):

There are two things that can happen. Either he draws a one and then a five, or he draws a five and then a one.

OpenStudy (anonymous):

So we have to test two things out. The chance of him getting a one then five and the chance of him getting a five then one.

OpenStudy (anonymous):

um kind of lost :/

OpenStudy (anonymous):

\[\frac{ 4 }{ 9 }*\frac{3}{8}\] will get you the probability of a one then five\[\frac{3}{9}*\frac{4}{8}\]will get you the probability of a five and then one.

OpenStudy (anonymous):

so basically what i said earlier

OpenStudy (anonymous):

ha XD

OpenStudy (anonymous):

Yeah, somewhat. The problem I'm thinking about is whether or not we're supposed to add these two probabilities. Since the question wants whether he gets a one and five, but doesn't give an order, it could be either of these.

OpenStudy (anonymous):

so will it be 0.16

OpenStudy (anonymous):

on the first one

OpenStudy (anonymous):

No need to convert to decimal form. Let's keep it in fraction form please.

OpenStudy (anonymous):

okay so will it be 1/6

OpenStudy (anonymous):

on both

OpenStudy (anonymous):

Yup.

OpenStudy (anonymous):

okay is that it or not

OpenStudy (anonymous):

thanks for your help guys

OpenStudy (anonymous):

@puppylife101 I did some testing to confirm my hypothesis. You ARE supposed to add the two probabilities.